# Theory and Interpretability of Quantum Extreme Learning Machines: a Pauli-transfer matrix approach

### v1

@@ V.2 Numerical results
We use here a complete measurement operator set of $4^{n}$ Pauli observables. In this setup, the reservoir unitary becomes immaterial [see Eq. 29], except for possible numerical conditioning issues. We generate trajectories using a 5th-order Runge-Kutta integrator with a time step of $\Delta t=0.01$. Reported values of the training error and forecast horizon are averaged over multiple initial conditions. After training the QELM, we obtain a classical surrogate model $\hat{\mathbf{f}}(\mathbf{u})$ by inserting into Eq. 25 the learned weights $\hat{\mathbf{w}}$.
@@
$\epsilon_{\text{train}}=\Big(\frac{1}{P}\sum_{t=1}^{P}\|\mathbf{u}_{t+1}-\hat{\mathbf{f}}(\mathbf{u}_{t})\|^{2}\Big)^{1/2},$ (59)
measures one-step accuracy of the trained predictor $\hat{\mathbf{f}}$. As a measure of multi-step prediction accuracy, the autonomously generated trajectory $\hat{\mathbf{u}}_{t+1}=\hat{\mathbf{f}}(\hat{\mathbf{u}}_{t})$ is compared to the true trajectory via the forecast horizon,
$T_{\text{fch}}=\inf\left\{t:\|\mathbf{u}_{t}-\hat{\mathbf{u}}_{t}\|>\sigma\right\}$ (60)
@@
Figure 9: QELM vs. its classical representation: Forecast obtained from a QELM [Eq. 1, ‘QRC’] with amplitude encoding (8) and $n=3$ qubits, trained on the rescaled Lorenz-63 model. This is macthed exactly by the forecast of the iterated classical representation [Eq. 25, ‘Surrogate’]. Both eventually diverge from the true trajectory due to the Lyapunov instability. Ground-truth trajectories are generated with a resolution of $\Delta t=0.01$. Time is given in model time units. For the scaling transformation in Eq. 57, we use parameters $\boldsymbol{\alpha}=0.004\cdot[1,1,1]$ and $\mathbf{m}=[-30,-30,-5]$.
@@
Table 2: Training RMS errors $\epsilon_{\text{train}}$ [Eq. 59] and forecast horizons $T_{\text{fch}}$ [Eq. 60] for the benchmark systems and (unitary-free) QELM setups discussed in the text. We generally use $n=3$ input qubits (no delay states) and a time resolution of $\Delta t=0.01$. $T_{\text{fch}}$ with its standard deviation (not error of the mean) is given in units of the Lyapunov time $T_{L}$. The column ‘type’ indicates whether we use the standard [Eq. 49] or rescaled dynamical system. In the latter case, $\alpha$ gives the rescaling factor in Eqs. 57 and 58 (same for all coordinates).
@@
Figure 10: Prediction performance (forecast horizon and RMS training error) under temporal multiplexing of a QELM trained on the Halvorsen model (58). We use amplitude encoding (9) with $n=3$ qubits, single-qubit Pauli-$Z$ observables, and (a) TFIM Hamiltonian (5a), (b) random Hamiltonian. The top horizontal axis shows the numerically obtained rank of the readout matrix $G$ [Eq. 28] relative to the maximum achievable rank $q=3^{n}$ [Eq. 30]. Forecast horizon is given in units of the Lyapunov time $T_{L}$; no additional features are added to the readout.
When employing temporal multiplexing with a QELM, the rank of the readout matrix $G$ [Eqs. 28 and 30] can serve as an indicator of the training accuracy, as illustrated in Fig. 10. This rank is upper bounded by the number of nonzero Pauli features q{q}, which, for the encoding (9) used here, is ${q}=3^{n}$. When $\mathrm{rank}(G)\simeq{q}$, all Pauli features are available at the readout. This can, in principle, improve performance when conditioning is controlled and exponential concentration effects can be avoided [51, 45]. Indeed, the QELM in Fig. 10(b) achieves training error and forecast horizon essentially identical to the ones reported in Table 2 obtained with a complete Pauli observable set. On the other hand, as seen in Fig. 10(a), performance is poor when the readout utilizes only a few effective observables. In the present setting, this is a consequence of using a structured Hamiltonian like the TFIM [Eq. 5a], which spreads observables less effectively over the full Hilbert space (see Fig. 4), such that $\mathrm{rank}(G)$ is capped by the rank of the PTM $R$.
@@ V.3 Interpretability
$\boldsymbol{\upphi}(\mathbf{u})=K\mathcal{B}(\mathbf{u}),$ (61)
@@
For the dynamical systems considered here, we can simply take $\mathcal{B}=\mathcal{B}_{r}$ to be the set of monomials in $u_{j}$ up to order $r$, of which there are $N={D+r\choose r}$. Accordingly, $K$ follows from Taylor-expanding $\boldsymbol{\upphi}(\mathbf{u})$ around some operating point $\bar{\mathbf{u}}$ (usually the data mean). As shown in Section IV.4.2, low-order monomials are indeed well constructible from $\boldsymbol{\upphi}$. However, due to the tensor product structure of $\boldsymbol{\upphi}$ [see Eq. 22] and the fact, that the derivatives of the single-qubit feature vector lie in a 2-dimensional space (vanishing Pauli-$Y$), $K$ typically has not full column rank when $N<q$. A consequence of this is that the expression of $\boldsymbol{\upphi}$ in terms of $\mathcal{B}$ is not unique. Nevertheless, deviations are small for the settings considered here  [^7] Deviations are small for $N<q$: in practice, for $n=D=3$ (thus $q=27$) and monomials of order $r=3$ (thus $N=19$ excluding the constant term), we get $\mathrm{rank}(K)=16$.
Figure 11 shows the learned flow maps, as approximated in Eq. 62, resulting from training on the (rescaled) Lorenz-63 and Halvorsen models (no feature engineering in the readout). Comparison to the theoretical expectation [Eq. 47] reveals good agreement especially for low-order coefficients. By contrast, deviations can be significant for higher-order coefficients. However, since they are multiplied by increasing powers of $\Delta t$, their influence on the overall training error is small.
@@ VI Summary
We have described a memoryless quantum reservoir computer (quantum extreme learning machine, QELM) in the Pauli transfer matrix (PTM) formalism, building upon [33, 25, 43, 46]. Writing states and channels in the Pauli basis (or any other orthogonal operator basis) turns the QELM into a sequence of real linear maps $T_{\mathcal{E}_{j}}$ acting on the vector of Pauli features $\boldsymbol{\upphi}(\mathbf{u})$ representing the encoded state $\rho_{\text{enc}}(\mathbf{u})$. These features carry the full dependence of the QELM on the encoded input $\mathbf{u}$ and determine the nonlinearities that the model can possibly express. Unitaries appear as orthogonal rotations on the traceless sector of $\boldsymbol{\upphi}(\mathbf{u})$, while general CPTP maps act affinely and typically contract the feature vector. Training a QELM reduces to classical linear regression over the components $\phi_{k}(\mathbf{u})$ and can be framed as a decoding problem of features mixed by quantum channels.
@@
While the PTM formalism has been previously been applied to QRCs [33, 25, 43, 46], we focused here on its implications for time-series prediction and model identification with QELMs. First, this approach allows one to cast the optimization of a QELM as an inverse problem, in which the desired target features imply an ideal encoding scheme and decodability from (preferably low-weight) Pauli features. Second, the PTM formalism provides a classical representation of the QELM [93, 94, 95, 47] and thereby renders it directly interpretable as a nonlinear vector (auto-)regression model (or next-generation RC [64]). These insights can help to construct reduced-order models in cases where the exact data-generating process is unknown and opens the door to further studies of black-box model identification [83, 96]. We also discussed the problem of temporal multiplexing and operator spreading [1, 33, 44, 59, 60] within the PTM framework, which can enable a more precise control over these and similar measurement enhancement strategies. Using the Pauli basis is justified by the measurement process, but not the only possibility. For future work, it could thus be interesting to study other orthogonal operator sets and assess the generality of the present findings.
@@
Figure 12: Pauli weight average (a) of Heisenberg-evolved measurement observables can predict the trainig error (b) and forecast horizon (c) of a QELM. We use a TFIM Hamiltonian (varying coupling parameters $J,h$, fixed $t=6$) with $n=3$ qubits and amplitude encoding (8) (no temporal multiplexing). The Pauli weight average is data-independent, while the QELM in (b,c) is trained on the Lorenz-63 model and employs 1-qubit Pauli-$Z$ observables in the readout, enhanced by monomials up to order 2 in order to increase sensitivity to nonlinear terms.
@@
It has previously been observed [97, 98, 99, 100, 101, 102, 103] that enhancing the quantum nature of a QML model often does not improve its performance on classical datasets. The present study adds further explanations of this finding from the viewpoint of learning dynamical systems with QELMs: first, a reservoir unitary mixes features and reversing this potentially detrimental effect requires complex measurement schemes (either high-order Pauli observables or temporal multiplexing), which, in turn, puts more demand on the linear decoder. We can demonstrate this mechanism directly by comparing the Pauli weight average of single-qubit observables to its training accuracy. As illustrated in Fig. 12 for a minimal QELM without temporal multiplexing, large Pauli weight average strongly correlates with high training error and short forecast horizon. In conventional QML models, instead, controlling this potentially detrimental feature mixing must be achieved by expensive training of variational gates. Second, optimal training error scaling requires learning of the flow map by a structurally matching feature library [88, 84]. Feature libraries provided by usual encoding schemes do not achieve this a priori and lead to surrogate models with significant errors (cf. Fig. 11). The forecast horizon $T_{\text{fch}}$ can be understood here akin to generalization error for static learning tasks: a nonzero $T_{\text{fch}}$ is only possible when the model has learned the actual rule governing the data  [^8] This is in contrast to the one-step training error, which can be decreased even by linear delay features.. Beside optimizing the measurement process [29], the present work thus points to the encoding as a central turning knob for performance of QELMs and, more generally, QML models (see, e.g., [105, 106] for insights along this line). How this surrogate-model argument gets modified when training on datasets without clear ground truths is an interesting avenue for future work [107].
@@
The $4^{n}$-dimensional space of Pauli features  [^9] Due to symmetries, the actual effective dimension is lower, and typically $\sim 3^{n}$ underlies the exponential frequency content found for QML models that employ rotational encoding [109, 49]. These and similar properties stemming from the exponential Hilbert space dimensionality point to a possible quantum advantages, but require specific measurement strategies. Moreover, simply increasing qubit number $n$ does not generally improve QRC performance, as overfitting and exponential concentration effects can become prevalent [51, 12, 45, 47]. In that regime, more sophisticated reservoir design strategies will be required to steer feature propagation through the quantum layers.
@@
$\mathbf{x}$ Input data ($\mathbf{x}\in\mathbb{R}^{D}$)
@@
$P_{k}$ $n$-qubit Pauli operators ($k=1,\ldots,d^{2}$)
@@
Table 3: Summary of notation used in this work.
@@ Appendix A Kernel representation
Training over a complete basis of measurement operators $\{M_{k}\}_{k=1}^{d}$ results in a quantum kernel form of the QELM [29]:
$f(\mathbf{x})=\sum_{j=1}^{P}\alpha_{j}^{*}K(\mathbf{x},\mathbf{x}_{j})=\mathrm{tr}(M^{*}\rho(\mathbf{x}))$ (63)
with coefficients $\alpha_{j}$ and the kernel [28]
$K(\mathbf{x},\mathbf{x}^{\prime})=\mathrm{tr}\big(\rho(\mathbf{x})\rho(\mathbf{x}^{\prime})\big)=\mathrm{tr}\big(\rho_{\text{enc}}(\mathbf{x})\rho_{\text{enc}}(\mathbf{x}^{\prime})\big)$ (64)
Notably, any data-independent unitary such as $U_{R}$ in (4) drops out from Eq. 63, implying that such a QELM is only dependent on the encoding state $\rho_{\text{enc}}(\mathbf{x})$. Furthermore, Eq. 63 allows to identify an optimal measurement operator:
@@
For the least-squares loss, the optimal parameters have the explicit representation $\boldsymbol{\alpha}^{*}=(\mathbf{K}+\lambda\mathbbm{1})^{-1}\mathbf{y}$, with the Gram matrix $K_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$.

### v2

@@ V.2 Numerical results
We use here a complete measurement operator set of $4^{n}$ Pauli observables. In this setup, the reservoir unitary becomes immaterial [see Eq. 29]. We generate trajectories using a 5th-order Runge-Kutta integrator with a time step of $\Delta t=0.01$. Reported values of the training error and forecast horizon are averaged over multiple initial conditions. After training the QELM, we obtain a classical surrogate model $\hat{\mathbf{f}}(\mathbf{u})$ by inserting into Eq. 25 the learned weights $\hat{\mathbf{w}}$.
@@
$\epsilon_{\text{train}}=\Big(\frac{1}{P}\sum_{t=1}^{P}\|\mathbf{u}_{t+1}-\hat{\mathbf{f}}(\mathbf{u}_{t})\|^{2}\Big)^{1/2}$ (55)
measures one-step accuracy of the trained predictor $\hat{\mathbf{f}}$. As a measure of multi-step prediction accuracy, the autonomously generated trajectory $\hat{\mathbf{u}}_{t+1}=\hat{\mathbf{f}}(\hat{\mathbf{u}}_{t})$ is compared to the true trajectory via the forecast horizon
$T_{\text{fch}}/\Delta t=\inf\left\{t:\|\mathbf{u}_{t}-\hat{\mathbf{u}}_{t}\|>\sigma\right\},$ (56)
@@
Figure 9: QELM vs. its classical representation: The forecast obtained from a QELM [Eq. 1] trained on the rescaled Lorenz-63 model is matched exactly by the forecast of its iterated classical (surrogate) representation [Eq. 25]. Both eventually diverge from the true trajectory due to the Lyapunov instability. The QELM employs amplitude encoding (8) with $n=3$ qubits. Ground-truth trajectories are generated with a resolution of $\Delta t=0.01$. Time is given in model time units. For the scaling transformation in Eq. 53, we use parameters $\alpha_{\kappa}=0.004$ and $\mathbf{m}=[-30,-30,-5]$.
@@
Table 2: Training RMS errors $\epsilon_{\text{train}}$ [Eq. 55] and forecast horizons $T_{\text{fch}}$ [Eq. 56] for the benchmark systems and (unitary-free) QELM setups discussed in the text. We generally use $n=3$ input qubits (no delay states) and a time resolution of $\Delta t=0.01$. $T_{\text{fch}}$ with its standard deviation (not error of the mean) is given in units of the Lyapunov time $T_{L}$. The column ‘type’ indicates whether we use the standard [Eq. 51] or rescaled dynamical system. In the latter case, $\alpha$ gives the rescaling factor in Eqs. 53 and 54 (same for all coordinates).
@@
Figure 10: Prediction performance (forecast horizon and RMS training error) under temporal multiplexing of a QELM trained on the Halvorsen model (54). We use amplitude encoding (9) with $n=3$ qubits, single-qubit Pauli-$Z$ observables, and (a) the TFIM Hamiltonian $H_{\text{TFIM}}^{\text{zz-x}}$ [Eq. 5a] or (b) a random Hamiltonian. The top horizontal axis shows the numerically obtained rank of the readout matrix $G$ [Eq. 28] relative to the maximum achievable rank $q=3^{n}$ [Eq. 30]. Forecast horizon is given in units of the Lyapunov time $T_{L}$; no additional features are added to the readout. For the TFIM variant $H_{\text{TFIM}}^{\text{xx-z}}$ [Eq. 5b], we obtain a maximum rank of $\mathrm{rank}(G)/3^{n}\simeq 0.22$, RMSE $\epsilon_{\text{train}}\simeq 7\times 10^{-4}$, and forecast horizon $T_{\text{fch}}\simeq 0$.
When employing temporal multiplexing, the rank of the readout matrix $G$ [Eq. 28] can serve as an indicator of the training accuracy, as illustrated in Fig. 10. This rank is upper bounded by the number of nonzero Pauli features q{q}, which, for the encoding (9) used here, is ${q}=3^{n}$. When conditioning is controlled and exponential concentration effects are avoided [53, 48], achieving $\mathrm{rank}(G)\simeq{q}$ makes essentially all Pauli features available to the readout and can thus improve performance. Indeed, the QELM in Fig. 10(b) achieves training error and forecast horizon essentially identical to the ones reported in Table 2 obtained with a complete Pauli observable set. By contrast, Fig. 10(a) shows that performance remains poor when the readout utilizes only a few effective observables. In the present setting, this is a consequence of using a structured Hamiltonian like the TFIM [Eq. 5a], which spreads observables less effectively over the full operator space, so that $\mathrm{rank}(G)\simeq\mathrm{rank}(R)\ll q$ [see Eqs. 31 and 4]. This effect is most pronounced when using the specific TFIM variant $H_{\text{TFIM}}^{\text{xx-z}}$ [Eq. 5b] with $Z$-type observables. Numerical experiments confirm the close relation (31) between the readout and PTM rank.
@@ V.3 Interpretability
[…]
$\boldsymbol{\upphi}(\mathbf{u})\simeq K\mathcal{B}(\mathbf{u}),$ (57)
@@
While there are several methods to determine $K$ (such as collocation), we perform here a Taylor expansion of $\boldsymbol{\upphi}(\mathbf{u})$ around some operating point $\bar{\mathbf{u}}$ (data mean). This is natural here, since, for the considered polynomial dynamical systems, we can take $\mathcal{B}=\mathcal{B}_{r}$ to be the set of monomials in $u_{j}$ up to order $r$, of which there are $N={D+r\choose r}$. Recall that low-order monomials are well constructible from the library $\boldsymbol{\upphi}$ [see Section IV.4.2]. However, due to the tensor product structure of $\boldsymbol{\upphi}$ [see Eq. 22] and the fact, that the derivatives of the single-qubit feature vector lie in a 2-dimensional space (vanishing Pauli-$Y$), $K$ typically has here not full column rank when $N<q$. A consequence of this is that only a $\mathrm{rank}(K)$-dimensional subspace of the library $\mathcal{B}$ is accessible through the encoding features $\boldsymbol{\upphi}$ and that hence some monomials in $\mathcal{B}$ are not independently identifiable. Nevertheless, deviations are small for the settings considered here  [^10] Deviations are small for $N<q$: in practice, for $n=D=3$ (thus $q=27$) and monomials of order $r=3$ (thus $N=19$ excluding the constant term), we get $\mathrm{rank}(K)=16$.
Figure 11 shows the learned flow maps, resulting from Eq. 58 after training a QELM with amplitude encoding [Eqs. 8 and 9] on the (rescaled) Lorenz-63 and Halvorsen models (no feature engineering in the readout). Comparison to the theoretical expectation [Eq. 49] reveals good agreement especially for low-order coefficients. While deviations can be large for higher-order coefficients, their influence on the overall training error is still small, since they are multiplied by increasing powers of $\Delta t$.
[…]
@@ VI Summary
We have employed the Pauli transfer matrix (PTM) formalism to describe a memoryless quantum reservoir computer (quantum extreme learning machine, QELM) with initial-state encoding and data-independent reservoir dynamics. Writing states and channels in the Pauli basis (or any other orthogonal operator basis) turns the QELM into a sequence of real linear maps $T_{\mathcal{E}_{j}}$ acting on the vector of Pauli features $\boldsymbol{\upphi}(\mathbf{u})$ representing the encoded state $\rho_{\text{enc}}(\mathbf{u})$. These features carry the full dependence of the QELM on the encoded input $\mathbf{u}$ and determine the nonlinearities that the model can possibly express. Unitaries appear as orthogonal rotations on the traceless sector of $\boldsymbol{\upphi}(\mathbf{u})$, while general CPTP maps act affinely. Training a QELM reduces to classical linear regression over the components $\phi_{k}(\mathbf{u})$ and can be framed as a decoding problem of features mixed by quantum channels.
@@
While the PTM formalism has been previously been applied to QRCs [36, 28, 46, 49], we focused here on its implications for time-series prediction and system identification with QELMs. First, this approach allows one to cast optimization of a QELM as an inverse problem: knowing the target features can point to an ideal encoding scheme as well as to constraints on the decodability paths. This goes beyond universal approximation theorems [26, 6, 27, 28, 29, 30] and allows one to gain more control over the training performance. Second, we have determined a classical and interpretable representation of the QELM in terms of a nonlinear vector (auto-)regression model (or “next-generation” RC [71]). The question of classical representations of QML models has been explored in numerous works [108, 109, 110, 50]. These insights can help to construct reduced-order models in cases where the exact data-generating process is unknown and address the problem of black-box system identification [98, 111]. Examining temporal multiplexing and operator spreading within the PTM framework demonstrated that greater reservoir randomness is advantageous under restricted measurement schemes. Using the Pauli basis is justified by the measurement process, but not the only possibility. For future work, it could thus be interesting to study other orthogonal operator sets and assess the generality of the present findings.
@@
Figure 13: Pauli weight average (a) of Heisenberg-evolved measurement observables can predict the training error (b) and forecast horizon (c) of a QELM. We use a TFIM Hamiltonian (varying coupling parameters $J,h$, fixed $t=6$) with $n=3$ qubits and amplitude encoding (8) (no temporal multiplexing). The Pauli weight average is data-independent, while the QELM in (b,c) is trained on the Lorenz-63 model and employs 1-qubit Pauli-$Z$ observables in the readout, enhanced by monomials up to degree 2 to increase sensitivity to nonlinear terms.
@@
Numerous previous studies noted that simply enhancing the quantum nature of a QML model often does not improve its performance on practical datasets [112, 113, 114, 115, 116, 117, 118]. The present study adds further explanations of this finding: first, a reservoir unitary mixes features and reversing this potentially detrimental effect requires complex measurement schemes (i.e., high-order Pauli observables, optimized pre-measurement unitaries, or temporal multiplexing schemes). We can demonstrate this mechanism directly by comparing the Pauli weight average of single-qubit observables to its training accuracy. As illustrated in Fig. 13, for such a minimal QELM without temporal multiplexing, large Pauli weight average strongly correlates with high training error and short forecast horizon on a time-series task. In conventional QML models, instead, feature mixing must be controlled by training of variational gates.
Second, learning of the flow map of a dynamical system benefits from a structurally aligned feature library [102, 105]. If the encoding scheme does not provide this, the resulting surrogate model can exhibit significant errors (cf. Fig. 11). In the case of chaotic systems, such errors are magnified by the Lyapunov instability and yield short forecast horizons $T_{\text{fch}}$  [^11] This is in contrast to the one-step training error, which can be decreased even by linear delay features.. Beside optimizing the measurement process [45, 32] and carefully structuring the Hamiltonian [43, 97, 50], the present work thus points to the encoding as a central turning knob for performance of QELMs and, more generally, QML models (see, e.g., [120, 121] for further insights along this line). How this surrogate-model argument gets modified when training on datasets with unknown ground truths is an interesting avenue for future work [122].
@@
The $4^{n}$-dimensional space of Pauli features  [^12] Due to symmetries, the actual effective dimension is lower, and typically $\sim 3^{n}$ can be compared to the exponential frequency content of QML models that employ gate-based rotational encoding [51, 124, 52]. These and similar properties stemming from the exponential Hilbert space dimensionality point to possible quantum advantages, but require specific measurement strategies. Moreover, simply increasing qubit number $n$ does not generally improve QRC performance, as ill-conditioning and exponential concentration effects can become prevalent [53, 12, 48, 50]. In that regime, more sophisticated reservoir design strategies will be required to steer feature propagation through the quantum layers.
@@
$\mathbf{x},\mathbf{u}$ Input data ($\mathbf{x},\mathbf{u}\in\mathbb{R}^{D}$)
@@
$P_{k}$ $n$-qubit Pauli operators ($k=0,\ldots,d^{2}-1$)
@@
Table 3: Summary of notation used in this work. A trained quantity is indicated by a hat (e.g., $\hat{\mathbf{w}}$).
@@ Appendix A Kernel representation
Training over a complete basis of measurement operators $\{M_{k}\}_{k=1}^{d^{2}}$ results in a quantum kernel form of the QELM [32]:
$f(\mathbf{x})=\sum_{j=1}^{P}\alpha_{j}^{*}K(\mathbf{x},\mathbf{x}_{j})=\mathrm{tr}(M^{*}\rho(\mathbf{x})),\qquad K(\mathbf{x},\mathbf{x}^{\prime})=\mathrm{tr}\big(\rho(\mathbf{x})\rho(\mathbf{x}^{\prime})\big)$ (59)
For the least-squares loss, the optimal parameters have the explicit representation $\boldsymbol{\alpha}^{*}=(\mathbf{K}+\lambda\mathbbm{1})^{-1}\mathbf{y}$, with the Gram matrix $K_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$ and the sample vector $\mathbf{y}\in\mathbb{R}^{P}$. If the reservoir consists only of a data-independent unitary, it drops out from Eq. 59,
[…]
showing that such a QELM only depends on the encoding state $\rho_{\text{enc}}(\mathbf{x})$. An optimal measurement operator can be identified from Eq. 59 as
@@
Compared to the equivalent primal formulation [Eq. 1], this representation is computationally more efficient when $P\ll d^{2}$.